# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ V-B Analytics techniques
With the development of FA, there’s a paramount requirement for effective scalable, and private techniques to aggregate insights across decentralized clients. It’s communication-intensive to upload encrypted data sets and it’s insecure to expose data even with noise while some transformation of data structure may effectively address these challenges. In this section, we discuss the data sketching and some interesting specialized data structures that artfully assist FA.
@@
Sketching. In FA, the pursuit of insight aggregation across decentralized data nodes demands techniques that are both efficient and scalable. Sketching stands out because it is a method designed with specific attributes tailored for the challenges posed by FA. These attributes [99] are query-specific (designed for specific insight extraction tasks), mergeable (allowing for easy aggregation from multiple clients), and extremely compact (ensuring minimal communication cost). Furthermore, various sketching methodologies have been developed to cater to specific data types and tasks. For instance, for set data, the Bloom filter is suitably employed for set intersection tasks. Similarly, the count sketch method is adept at identifying heavy hitters. To address the paramount privacy requirements in FA, sketching often collaborates with privatization methods such as noise addition, randomized responses, and hash functions to ensure the privacy and security of the aggregated insights. Together, these techniques present a robust and private solution for insight aggregation in FA.
@@
Bloom filters [100] is a space-efficient probabilistic data structure used to test whether an element belongs to a set with possible false positives and no false negatives. Bloom filter and its variants are first introduced to federated analytics in Private Set Intersection tasks [101, 102, 103, 104]. Researchers in [101] combined Bloom filters with Goldwasser Micali homomorphic encryption to compute the intersection which is secure in malicious model. In [102], authors proposed a (PSI) method using Bloom filters with a secure multiplication protocol SEPIA to obtain an intersection of Bloom filters for each party to find their intersections. Authors in [103] reduced the hash operations to propose a faster protocol for the PSI task by integrating oblivious transfer with a garbled Bloom filter (GBF). The essential difference between the GBF and Bloom filters is that the GBF uses an array of $\lambda$-bit string where $\lambda$ is a security parameter while Bloom filters use an array of bits. Specifically, the client computes its sets to a Bloom filter and the server also computes its set to a GBF. Then an oblivious transfer protocol is conducted to give the client a GBF of intersection which is finally queried by the client to obtain the result. There were continuing works on Dong et al. [103] such as optimization of performance based on random OT extension[104], enhancement in its malicious-secure variant [105], and multi-parties intersection[96, 106]. RAPPOR [15] introduced by Google is a widely applicable and practical data collection mechanism providing strong privacy guarantees with high utility. The basic idea of RAPPOR is applying randomized responses to Bloom filters that each client reports each bit of her Bloom filter with possible untruthful responses to the server for privacy guarantees.
@@
TABLE V: Summary of sketching.
Reference Task Sketching method Cooperated privatization technique
Kerschbaum [101] PSI between 2 parties Bloom filters Goldwasser Micali Encryption
Many et al. [102] Multi-parties PSI Bloom filters SEPIA library
GBF [103] PSI between 2 parties Bloom filters Oblivious Transfer
Pinkas et al. [104] PSI between 2 parties Bloom filters Oblivious Transfer
Rindal et al. [105] PSI between 2 parties Bloom filters Oblivious Transfer&Choose-and-cut
Kolesnikov et al. [96] Multi-parties PSI Bloom filters Oblivious Transfer&Pseudorandom functions
Inbar et al. [106] Multi-parties PSI Bloom filters Oblivious Transfer
RAPPOR [15] Data federation Bloom filters Randomized responses
CMS&HCMS [82] Dictionary-related tasks Count sketching Hash functions&Hadamard transform
TreeHist [25] Heavy hitters Count sketching Hash functions&Hadamard transform
HH&PEM [44] Heavy hitters Count sketching Noise addition
@@
Specialized data structures. In the context of FA where data analysis is performed across federated devices and servers, some specialized data structures are proposed to help the server to derive the analytics results on particular tasks. In [18], authors proposed the TrieHH algorithm discovering frequently typed words in edge devices, which is an iterative algorithm to find heavy hitters with the Prefix Tree. The Prefix tree, or the Trie, is a tree-like data structure used to store strings like words. Each node in the Trie represents a single character of a string and all end nodes represent an end identifier so that each end node stands for a string and all sibling nodes own a common prefix. In their TrieHH algorithm, a subset of clients is sampled in each iteration and the Trie is updated based on their data points after filtering by a threshold until convergence. This sampling-and-threshold algorithm is used to provide central differential privacy instead of the usual noise addition. There’s a following work of it, TrieHH++[43], which answered more general queries that not only identified heavy hitters but also provided their estimated frequency. In [107], researchers proposed OptPrefixTree based on previous work. They suggest a flexible algorithm to increase efficacy by utilizing adaptive segmentation, intelligent data selection, and deny lists.
@@
There’s also a tree-based method utilized in [29] to guarantee robustness and prevent the aggregator from cheating, which is a summation tree. Authors proposed the Honeycrisp system which uses sparse vector theorem to schedule the DP budget. The aggregator only has ability to conduct summation on collected encrypted data. The committee composed of some clients can decrypt the result and compare it with the guess from the aggregator to decide whether to reveal it. When conducting the total summation in the aggregation phase, the aggregator is asked to generate a summation tree in which each parent node owns two children nodes and is the sum of them under homomorphic encryption. The tree is then checked by clients with each client verifying one pair of a parent and its children to prevent the aggregator from manipulating the aggregated result.
@@ V-C Deployment and optimization techniques
Beyond the basic privacy afforded by restricting raw data within the data owner, there’s the complex challenge of deploying effective techniques for optimization in some scenarios. These techniques are essential to navigate the distinct problem FA presents in terms of computation, communication, data heterogeneity, incentive mechanism, utility, and privacy. In this section, we list some techniques concentrating on the deployment and optimization problems in FA.
@@
Optimization theory. The optimization challenges in federated analytics differ significantly from those in a centralized setting. As mentioned in [108], federated optimization mainly focuses on communication efficiency, data heterogeneity, computation, and privacy constraints.
@@
In [109], FedVision is introduced, optimizing resource usage in video analytics. This system minimizes network resource consumption and maximizes computational efficiency by integrating black-box optimization with Neural Processes, tailored for dynamic network conditions. The study in [110] presents Edge-DemLearn, an approach that maximizes learning performance and system efficiency. By leveraging distributed computing infrastructure, Edge-DemLearn is a sub-optimal two-sided many-to-one matching algorithm optimizing both the allocation of resources and the generalization of models. This method effectively overcomes the inherent limitations of traditional Federated Learning by enhancing model generalization and optimizing the management of user equipment associations within a distributed learning framework. Paper [111] represents a method that combines federated analytics with 5G technology to optimize data collection and processing in vehicular networks. This approach aims to minimize latency by employing a federated analytics-based solution and a D/M/1 queuing systems-based mathematical model to analyze end-to-end latency. They utilize numerical optimization methods like linear approximation (COBYLA), the Trust Region Constrained Algorithm (TRCA), and the Sequential Least SQuares Programming (SLSQP) to optimize system parameters, focusing on reducing waiting times and maximizing information value. S. Shi et al. in [112] introduced a Distributionally Robust Optimization (DRO) paradigm to optimize robustness in network traffic classifier learning under noisy labels. DRO is a modeling approach that makes decisions under uncertainty by minimizing the worst-case cost across all possible distributions in a constructed uncertainty set. This strategy maximizes classifier accuracy despite the variability in data quality.
@@
Game theory and incentive design. In federated settings, incentive mechanisms are crucial for motivating participants to contribute their private data, communicational resources, and computational resources. Incentive design typically balances individual rewards with collective goals to guarantee that participants are fairly compensated and to maintain the efficiency of the system. By employing game theory approaches such as contract theory and multi-leader-follower games, these incentives aim to align participant actions with the overall objectives of the federated tasks. The structure of incentive mechanisms in FA is considerably similar to those in FL, where collaborative data processing and performing tasks require nearly the same strategies. Game theory in federated analytics emerges as a critical tool for resolving complex interactions and optimizing cooperative strategies across various domains. It efficiently addresses challenges in privacy and incentivizes participation, scalability, and effectiveness in decentralized systems.
@@
In [113], researchers introduce CROWDFA which employs federated analytics in mobile crowdsensing and balances data aggregation, incentive design, and privacy preservation. It uses additive secret sharing for privacy and a novel incentive mechanism, PRAED, to encourage participants’ involvement with privacy protection. Authors in [114] integrate blockchain with federated learning in healthcare metaverses, focusing on a user-centric incentive mechanism. It introduces AoI as a metric for data freshness and employs contract theory to incentivize data sharing. The framework uses Prospect Theory to address the service provider’s decision-making under uncertainty, aiming to enhance immersive healthcare experiences by ensuring data freshness and privacy. Another incentive mechanism is developed in [115] for federated learning and analytics systems with multiple tasks. It uses a multi-leader-follower game where rewards are set to motivate data owners, who in turn decide their participation level. The paper provides algorithms for optimal strategy formulation, ensuring effective resource allocation and addressing the challenges of multi-task federated systems. S. Shi et al. in [116] focuses on updating HD maps for autonomous driving using a game-theoretical model. It proposes an overlapping coalition formation game where vehicles collaborate in coalitions to update map data for maximizing utility and map quality. It addresses privacy and incentivizes participation by allowing vehicles to join multiple coalitions, which benefits the scalability and accuracy of map updates. In [117], researchers cope with privacy and pricing in data markets. It proposes a federation model where data providers form coalitions with differential privacy. They introduce a method for determining collective data prices based on privacy levels and employ the Shapley value from game theory for fair earnings distribution, which benefits both data providers and consumers by balancing privacy concerns and financial incentives.
@@
Sampling. To aggregate insights across these distributed clients, federated analytics often utilizes sampling techniques in terms of enhancing the privacy and improving the performance. There’s also an optimization-wise problem for sampling techniques and the accuracy as shown in figure 5. 1) When sampling is utilized, the total amount of data decreases so that the privacy budgets are amplified. The accuracy of results is improved since the federated data are less perturbed with more privacy budget. 2) While intuitively, less data analyzed leads to more variance of the results that sampling may cause accuracy reduction.
@@
In early research within the domain of differential privacy, the sampling technique is referenced and employed for the purpose of privacy amplification[118]. Instead of examining the full data from all clients, using either data from a sample of clients or each client’s sampled subset of data can enhance privacy protection in terms of amplify the privacy bounds of a differentially private mechanism under various conditions. In other words, the utilization of sampling has ability to improve the accuracy with the ”amplified” privacy budgets. In most time, sampling is combined with other privatization techniques like the noise injection to further protect data. Researchers in [38] proposed the LDPMiner, a two phases mechanism for heavy hitters task which applied sampled randomizer (combining sampling with privatization techniques like RAPPOR and Succinct Histogram) to each phase. They conducted the experiments on synthetic data and one of results is the accuracy of targets from sampled randomizer outperforms that from naive randomizer under the same privacy budgets. It’s worth noting that there’s a born-in limitation of sampling that it’s always require large enough total data to ensure its effectiveness, which is also indicated by another experimental results in [38]. There’s a continuing work on LDPMiner, which proposed SVIM [41], a protocol for searching frequent items under set-valued LDP setting. They attempted new cooperated privatization techniques with sampling and found that some techniques benefit privacy amplification from sampling like GRR (Generalized Random Response) while some doesn’t such as OLH (Optimized Local Hashing). They explained the interesting phenomenon that the relation between reported and input values needs to satisfy a ”many-to-one-wise” mapping for benefit of privacy amplification from sampling, i.e., reported value after encryption ”support” one input value instead of multiple values [41]. Also, sampling can improve efficiency of system attribute to the reduction in amount of data transmitted and processed while facing with the concern of accuracy loss. In [49], authors proposed a private data federation system SAQE (Secure Approximate Query Evaluator) and denoted that sometimes sampling improves not only the efficiency but also the accuracy, i.e. utilizing less data in the computation leads to improvement in accuracy. Given invariant privacy budgets, the sampling error decreases as the sampling rate grows up, which means it’s required more noise injection to maintain the differential privacy. Next, the total error summed by sampling error and error from noise is a convex function of sampling rate for some tasks, meaning that sampling with proper rate contributes to an increasing accuracy. Now that sampling has the ability to introduce error to guarantee privacy, there are also attempts on merely applying sampling instead of as a bonus privacy amplifier with other techniques like noise injection.
TABLE VI: Summary of sampling.
Reference Sampled Object Sampling rate Sampling method Cooperated privatization technique
LDPMiner [38] Data of each client 1 over data number for each client Uniform Succint Histogram&RAPPOR
SVIM [41] Data of each client 1 over data number for each client Uniform GRR&OLH
SAQE [49] Data of each client Optimized result based on the task Bernoulli Noise addition

### v3

@@ V-B Analytics techniques
With the development of FA, there’s a paramount requirement for effective and accurate techniques to aggregate insights across decentralized clients. It is communication-intensive to upload encrypted datasets and it is insecure to expose data even with noise while some transformation of data structure may effectively address these challenges. In this section, we discuss the data sketching and some interesting specialized data structures that artfully assist FA.
V-B1 Sketching
@@
Data sketching is a technique for creating compact summaries (sketches) of large datasets that approximate key properties, such as cardinality, frequency, or quantiles [171, 172, 173], enabling efficient computation with reduced memory and processing requirements. In FA, the pursuit of insight aggregation across decentralized data nodes demands techniques that are both efficient and scalable. Sketching stands out because it is a method designed with specific attributes tailored for the challenges posed by FA. These attributes [174] are query-specific (designed for specific insight extraction tasks), mergeable (allowing for easy aggregation from multiple clients), and extremely compact (ensuring minimal communication cost). Furthermore, various sketching methodologies have been developed to cater to specific data types and tasks. For instance, for set data, the Bloom filter is suitably employed for set intersection tasks. Similarly, the count sketch method is adept at identifying heavy hitters. To address the paramount privacy requirements in FA, sketching often collaborates with privatization methods such as noise addition, randomized responses, and hash functions to ensure the privacy and security of the aggregated insights. Together, these techniques present a robust and private solution for insight aggregation in FA. Unlike FL, which primarily focuses on model training, FA leverages sketching to handle specific data aggregation tasks such as set intersections and heavy hitter identification. Moreover, sketching in FA often integrates privacy-preserving techniques to ensure the confidentiality of the aggregated insights. These attributes highlight the unique and critical role of sketching in FA, distinguishing its application from other areas.
@@
Bloom filters [175] is a space-efficient probabilistic data structure used to test whether an element belongs to a set with possible false positives and no false negatives. Bloom filter and its variants are first introduced to FA in Private Set Intersection tasks [176, 177, 178, 179]. Researchers in [176] combined Bloom filters with Goldwasser Micali HE to compute the intersection which is secure in a malicious model. In [177], authors proposed a PSI method using Bloom filters with a secure multiplication protocol SEPIA to obtain an intersection of Bloom filters for each party to find their intersections. Authors in [178] reduced the hash operations to propose a faster protocol for the PSI task by integrating oblivious transfer with a garbled Bloom filter (GBF). The essential difference between the GBF and Bloom filters is that the GBF uses an array of $\lambda$-bit string where $\lambda$ is a security parameter while Bloom filters use an array of bits. Specifically, the client computes its sets to a Bloom filter and the server also computes its set to a GBF. Then an oblivious transfer protocol is conducted to give the client a GBF of intersection which is finally queried by the client to obtain the result. There were continuing works on Dong et al. [178] such as optimization of performance based on random OT extension[179], enhancement in its malicious-secure variant [180], and multi-parties intersection[168, 181]. RAPPOR [25] introduced by Google is a widely applicable and practical data collection mechanism providing strong privacy guarantees with high utility. The basic idea of RAPPOR is applying randomized responses to Bloom filters so that each client reports each bit of her Bloom filter with possible untruthful responses to the server for privacy guarantees.
@@
Lessons learned. Sketching in FA highlights its critical role in addressing scalability, efficiency, and privacy challenges. Sketching techniques, such as Bloom filters and Count Sketch, stand out for their compactness, mergeability, and task-specific designs, making them well-suited for FA tasks like heavy hitter identification and set intersections. These methods ensure low communication costs while enabling accurate insight aggregation from distributed nodes. To preserve privacy, sketching is often integrated with techniques like noise addition, randomized responses, and hash functions. For example, Bloom filters are enhanced with differential privacy mechanisms like randomized responses in tools such as RAPPOR, while Count Sketch is combined with Hadamard transforms for tasks like frequency estimation. These innovations underscore the adaptability of sketching in balancing privacy and utility in decentralized environments. Sketching proves indispensable for scalable, privacy-preserving insight aggregation in FA.
V-B2 Specialized data structures
@@
In the context of FA where data analysis is performed across federated devices and servers, some specialized data structures are proposed to help the server to derive the analytics results on particular tasks. These specialized data structures are uniquely tailored to the requirements of FA, facilitating accurate and secure data aggregation. In [28], authors proposed the TrieHH algorithm discovering frequently typed words in edge devices, which is an iterative algorithm to find heavy hitters with the Prefix Tree. The Prefix tree, or the Trie, is a tree-like data structure used to store strings like words. Each node in the Trie represents a single character of a string and all end nodes represent an end identifier so that each end node stands for a string and all sibling nodes own a common prefix. In their TrieHH algorithm, a subset of clients is sampled in each iteration and the Trie is updated based on their data points after filtering by a threshold until convergence. This sampling-and-threshold algorithm is used to provide CDP instead of the usual noise addition. There’s a following work of it, TrieHH++[70], which answered more general queries that not only identified heavy hitters but also provided their estimated frequency. In [67], researchers proposed OptPrefixTree based on previous work. They suggest a flexible algorithm to increase efficacy by utilizing adaptive segmentation, intelligent data selection, and deny lists.
@@
There is also a tree-based method utilized in [37] to guarantee robustness and prevent the aggregator from cheating, which is a summation tree. Authors proposed the Honeycrisp system which uses sparse vector theorem to schedule the DP budget. The aggregator only can conduct summation on collected encrypted data. The committee composed of some clients can decrypt the result and compare it with the guess from the aggregator to decide whether to reveal it. When conducting the total summation in the aggregation phase, the aggregator is asked to generate a summation tree in which each parent node owns two children nodes and is the sum of them under HE. The tree is then checked by clients with each client verifying one pair of a parent and its children to prevent the aggregator from manipulating the aggregated result.
Lessons learned. Specialized data structures in FA, such as prefix trees (e.g., TrieHH and TrieHH++) and summation trees, enable efficient, accurate, and privacy-preserving data aggregation. Trie-based methods effectively identify heavy hitters with CDP using iterative updates, while optimized versions like OptPrefixTree enhance flexibility and efficiency through adaptive segmentation and intelligent data selection. Summation trees, as seen in Honeycrisp, ensure robust aggregation with HE and client-side verification, preventing manipulation. These tailored structures highlight the importance of combining privacy, robustness, and task-specific adaptability in FA systems.
@@ V-C Deployment and optimization techniques
Beyond the basic privacy afforded by restricting raw data within the data owner, there exists the complex challenge of deploying effective techniques for optimization in some scenarios. These techniques are essential to navigate the distinct problem FA presents in terms of computation, communication, data heterogeneity, incentive mechanism, utility, and privacy. In this section, we list some techniques concentrating on the deployment and optimization problems in FA.
V-C1 Optimization theory
@@
The optimization challenges in FA differ significantly from those in a centralized setting. As mentioned in [182], federated optimization mainly focuses on communication efficiency, data heterogeneity, computation, and privacy constraints.
@@
In [23], FedVision is introduced, optimizing resource usage in video analytics. This system minimizes network resource consumption and maximizes computational efficiency by integrating black-box optimization with Neural Processes, tailored for dynamic network conditions. The study in [127] presents Edge-DemLearn, an approach that maximizes learning performance and system efficiency. By leveraging distributed computing infrastructure, Edge-DemLearn is a sub-optimal two-sided many-to-one matching algorithm optimizing both the allocation of resources and the generalization of models. This method effectively overcomes the inherent limitations of traditional FL by enhancing model generalization and optimizing the management of user equipment associations within a distributed learning framework. [126] represents a method that combines FA with 5G technology to optimize data collection and processing in vehicular networks. This approach aims to minimize latency by employing an FA-based solution and a D/M/1 queuing systems-based mathematical model to analyze end-to-end latency. They utilize numerical optimization methods like linear approximation (COBYLA), the Trust Region Constrained Algorithm (TRCA), and the Sequential Least SQuares Programming (SLSQP) to optimize system parameters, focusing on reducing waiting times and maximizing information value. Shi et al. [128] introduced a Distributionally Robust Optimization (DRO) paradigm to optimize robustness in network traffic classifier learning under noisy labels. DRO is a modeling approach that makes decisions under uncertainty by minimizing the worst-case cost across all possible distributions in a constructed uncertainty set. This strategy maximizes classifier accuracy despite the variability in data quality.
V-C2 Game theory and incentive design
@@
In federated settings, incentive mechanisms are crucial for motivating participants to contribute their private data, communicational resources, and computational resources. Incentive design typically balances individual rewards with collective goals to guarantee that participants are fairly compensated and to maintain the efficiency of the system. By employing game theory [183] approaches such as contract theory and multi-leader-follower games, these incentives aim to align participant actions with the overall objectives of the federated tasks. The structure of incentive mechanisms in FA is considerably similar to those in FL [184], where collaborative data processing and performing tasks require nearly the same strategies. Game theory in FA emerges as a critical tool for resolving complex interactions and optimizing cooperative strategies across various domains. It efficiently addresses challenges in privacy and incentivizes participation, scalability, and effectiveness in decentralized systems.
@@
In [130], researchers introduce CROWDFA which employs FA in mobile crowdsensing and balances data aggregation, incentive design, and privacy preservation. It uses additive secret sharing for privacy and a novel incentive mechanism, PRAED, to encourage participants’ involvement with privacy protection. Authors in [185] integrate blockchain with FL in healthcare metaverses, focusing on a user-centric incentive mechanism. It introduces AoI as a metric for data freshness and employs contract theory to incentivize data sharing. The framework uses Prospect Theory to address the service provider’s decision-making under uncertainty, aiming to enhance immersive healthcare experiences by ensuring data freshness and privacy. Another incentive mechanism is developed in [131] for FL/FA systems with multiple tasks. It uses a multi-leader-follower game where rewards are set to motivate data owners, who in turn decide their participation level. The paper provides algorithms for optimal strategy formulation, ensuring effective resource allocation and addressing the challenges of multi-task federated systems. Shi et al. [129] focuses on updating HD maps for autonomous driving using a game-theoretical model. It proposes an overlapping coalition formation game where vehicles collaborate in coalitions to update map data for maximizing utility and map quality. It addresses privacy and incentivizes participation by allowing vehicles to join multiple coalitions, which benefits the scalability and accuracy of map updates. In [186], researchers cope with privacy and pricing in data markets. It proposes a federation model where data providers form coalitions with DP. They introduce a method for determining collective data prices based on privacy levels and employ the Shapley value from game theory for fair earnings distribution, which benefits both data providers and consumers by balancing privacy concerns and financial incentives.
Lessons learned. Incentive mechanisms are essential for motivating participants in FA, balancing individual rewards with collective goals while ensuring system efficiency. Game theory-based approaches, such as contract theory and multi-leader-follower games, effectively align participant actions with overall objectives, addressing challenges in privacy, scalability, and participation. Examples like CROWDFA and blockchain-integrated healthcare metaverses demonstrate how tailored incentive designs, such as additive secret sharing and privacy-preserving metrics, drive collaboration while maintaining privacy. Additionally, multi-task incentive mechanisms and coalition formation models for applications like HD map updates and data markets highlight the importance of fair resource allocation and financial rewards, often leveraging tools like the Shapley value to balance privacy and utility. These mechanisms showcase the critical role of game theory and innovative strategies in fostering effective and scalable FA systems.
@@ V-C3 Sampling
Data sampling is the process of selecting a subset of data points from a larger dataset to analyze or draw conclusions about the whole, often to reduce computational cost while maintaining statistical representativeness [187, 188]. To aggregate insights across these distributed clients, FA often utilizes sampling techniques in terms of enhancing privacy and improving performance. There’s also an optimization-wise problem for sampling techniques and the accuracy as shown in figure 7. 1) When sampling is utilized, the total amount of data decreases so that the privacy budgets are amplified. The accuracy of results is improved since the federated data are less perturbed with more privacy budget. 2) While intuitively, less data analyzed leads to more variance in the results where sampling may cause accuracy reduction.
@@
In early research within the domain of DP, the sampling technique is referenced and employed for privacy amplification[55]. Instead of examining the full data from all clients, using either data from a sample of clients or each client’s sampled subset of data can enhance privacy protection in terms of amplifying the privacy bounds of a differentially private mechanism under various conditions. In other words, the utilization of sampling can improve the accuracy with the ”amplified” privacy budgets. Most time, sampling is combined with other privatization techniques like noise injection to further protect data. Researchers in [62] proposed the LDPMiner, a two-phase mechanism for heavy hitters tasks which applied sampled randomizer (combining sampling with privatization techniques like RAPPOR and Succinct Histogram) to each phase. They conducted the experiments on synthetic data and one of the results is the accuracy of targets from sampled randomizers outperforms that from naive randomizers under the same privacy budgets. It is worth noting that there’s a built-in limitation of sampling that it always requires large enough total data to ensure its effectiveness, which is also indicated by another experimental result in [62]. There is continuing work on LDPMiner, which proposed SVIM [68], a protocol for searching frequent items under a set-valued LDP setting. They attempted new cooperated privatization techniques with sampling and found that some techniques benefit privacy amplification from sampling like Generalized Random Response while others do not such as Optimized Local Hashing. They explained the interesting phenomenon that the relation between reported and input values needs to satisfy a ”many-to-one-wise” mapping for the benefit of privacy amplification from sampling, i.e., reported value after encryption ”support” one input value instead of multiple values [68]. In addition, researchers utilize sampling in TrieHH [28], TrieHH++ [70], and OptPrefixTree [67] to choose a subset of clients in each iteration, where it is proved that sampling saves privacy budget.
However, sampling can improve the efficiency of the system attributed to the reduction in the amount of data processed while facing the concern of accuracy loss. In [77], authors proposed a private data federation system Secure Approximate Query Evaluator (SAQE) and denoted that sometimes sampling improves not only the efficiency but also the accuracy, i.e., utilizing fewer data in the computation leads to improvement in accuracy. Given invariant privacy budgets, the sampling error decreases as the sampling rate grows, which means it requires more noise injection to maintain DP. Next, the total error summed by sampling error and error from noise is a convex function of the sampling rate for some tasks, meaning that sampling with the proper rate contributes to increasing accuracy. This utilization of sampling to increase the accuracy quantitatively of aggregation results by amplification of the privacy budget is uniquely used in the FA task. Utilization of the sampling technique in FA is highly associated with our taxonomy of data owner type, where its utilization is different for cross-silo FA and cross-device FA. In cross-silo FA, sampling is conducted within the data points held by clients and obtains benefits of privacy amplification and reduction of overheads caused by cryptography. However, it requires the clients (silos) to be trusted to perform sampling correctly on their data. Meanwhile, sampling in cross-device FA is conducted as the server randomly samples the clients for participating in FA. Such sampling is the vanilla setting in most of the cross-device federated computation systems. However, it requires the central server to be trusted to some extent. It is challenging to maintain the secrecy of sampling in cross-device FA, i.e., nobody knows anybody is sampled, which is an important prerequisite of privacy amplification.
Lessons learned. Sampling techniques in FA are powerful tools for enhancing privacy, improving efficiency, and even increasing accuracy under certain conditions. By reducing the amount of data processed, sampling amplifies privacy budgets, allowing less perturbation and improved accuracy in DP mechanisms. However, the trade-off is a potential increase in variance due to fewer data points, highlighting the importance of balancing sampling rates to optimize accuracy and efficiency. Additionally, maintaining the secrecy of sampling in a one-data-per-person scenario—where no one knows who has been sampled—can be costly, making it less feasible for applications with strict privacy requirements.
In the previous sections, we have summarized many key factors of existing FA works and potential FA designs, including its general definition, taxonomy, challenges, and enabling techniques. In the following five sections, we introduce existing solutions on FA in detail. For reading convenience, we classify FA solutions based on their host data analytics tasks as mentioned in Section III-D1. Some sections will be further divided into subsections if the class of data analytics tasks can be further classified. Within each section/subsection, the existing FA solutions on a particular class of data analytics tasks will be introduced ordered by their privatization methodology.